OpenStudy (anonymous):

approximate using the trapezoidal rule by dividing the interval 1

OpenStudy (anonymous):

\[\int\limits_{1}^{5}6/x \]

OpenStudy (anonymous):

this is my approach but it appears wrong according to the answers: \[\int\limits_{1}^{5}6/x\] \[\approx\] \[\int\limits_{1}^{2}6/x +\int\limits_{2}^{3}6/x +\int\limits_{3}^{4}6/x +\int\limits_{4}^{5}6/x\] \[\approx\] 2(6+3)+2(3+2)+2(2+3/2)+2(6/5+3/2) \[\approx\] 40.4